OpenStudy (anonymous):

f=ma if a =0 is the equation valid or invalid....

OpenStudy (turingtest):

If a=0 that means net force F=0. (i.e. the forces cancel and the object does not accelerate) So yes, the equation is perfectly valid. Now if m=0... well that's another story :)

OpenStudy (anonymous):

the equation is valid....u re correct man

OpenStudy (anonymous):

the equation is valid. when a=0 (i.e v is not change ) \[\rightarrow\] f=0

OpenStudy (anonymous):

when a=0, u won't be needin this equation at all, because there won't be any force to measure or perceive

OpenStudy (turingtest):

@abhinayreddy That is certainly not true. Just because the net force is zero doesn't mean you can't use the equation. All a=0 means is that the object is in equilibrium, which is usually due to the fact that whatever forces are acting on the object are cancelling. For instance if we want to find the force of static friction on an object, we need to know that \[\sum F=F_{external}+F_g+N+f=0\](all of the above are vectors) Here the object is not accelerating (i.e. a=0) but the equation is completely necessary.

OpenStudy (anonymous):

@TuringTest what i was saying is that when a=0 you won't have any net force to perceive at all and other than to know the weight (which can be found by other means as well) how can we make use of this equation? and i agree whether a=0 or not, this equation is valid in newtonian physics.